# Text-Guided Variational Image Generation for Industrial Anomaly Detection and Segmentation

### v1

@@ 2 Related Works
Text-to-image Generation. Recently, text-to-image generation models [17, 14] are widely studied, which predict or generate the next pixel values based on the given text and all previous predictions of pixel values. For example, DALL-E [18] takes text input from GPT-3 to predict the next pixels, which are then added to the initial image to be placed as the subsequent input along with the text. By repeating the process multiple times, DALL-E can generate an entire image. Generative Adversarial Network (GAN) [10] is known for high-quality and realistic image generation without a prompt, and Contrastive Language-Image Pre-training (CLIP) is another neural network that can determine how well a caption or prompt matches an image. Through VQGAN-CLIP [5], where VQGAN [8] creates an image, CLIP [17] checks whether the generated image from VQGAN matches the text. When the process is repeated, the created image gradually resembles the input text.
@@ 3 Preliminary Analysis
We performed preliminary experiments with a simple scenario to show the effectiveness of a newly generated image method for anomaly detection tasks. In the scenario, we compare the performance between the baseline anomaly detection model and the same model using one additional training image searched on the web or generated by an image generator. We consider a single hazelnut image from the MVTecAD dataset [3], utilizing PatchCore [19] for the baseline anomaly detection model.
@@
The results of preliminary experiments are represented in Table 1, showing that performance is affected depending on the generated images. Interestingly, preliminary results show that good image quality does not necessarily help improve performance as presented in the low performance of DALL-E. Meanwhile, given the high performance of Web Image and Midjourney, we can enhance performance when elements such as the outline and angle of the original image are preserved in the additional image.
Table 1: Preliminary experiment using Patchcore. [19] Using just one image of hazelnut, this experiment examines the performance changes of anomaly detection models utilizing the additional training images retrieved from external webs or generated by generative models using text prompts. The results highlight the need for a well-designed generation model to boost anomaly detection performance by adding extra training images.
@@ 4 Methodology
As shown in Fig. 3, our framework comprises the keyword-to-prompt generator, variance-aware image generator, and text-guided knowledge integrator. The keyword-to-prompts generator produces a group of prompts consisting of different word combinations based on the essential words presented in the input text, then selects the best prompt that contains comparable information to the input image among the generated ones. The variance-aware image generator generates a set of non-defective images, encoding the visual features of the non-defective images into the normal distribution to preserve their variance. The generation process is iterated with the generator update, and the text-guided knowledge integrator determines the optimal set of generated images by assessing the latent distribution similarity between the image set and the text prompts generated by the keyword-to-prompt generator. Finally, the optimal set of generated images is utilized as the additional training set for the baseline anomaly detection model.
Figure 3: Overview of our framework. Our framework is composed of a text-guided knowledge integrator, a variance-aware image generator, and keyword-to-prompt generator. The keyword-prompt generator generates a group of prompts consisting of word combinations based on the input target name, and then selects the best prompt among the generated prompts that contains similar information to the input image. The variance-aware image generator encodes the visual features of a non-defective image as a normal distribution and embeds the variance of the encoded features. The text-guided knowledge integrator is responsible for encoding the generated images from the distributed recognition image generator, evaluating the similarity of the resulting features to the optimal prompts and deriving the optimal image. The generated non-defective images are added into the non-defective data pool to train the baseline anomaly detection models.
@@ 4.2 Text-guided Knowledge Integrator
Figure 4: Generalization test for anomaly detection in MVTecAD dataset. The first row shows the average improving score across different baselines and varying numbers of non-defective images. The second rows present the average score for the highest-improving five classes. Detailed all classes experimental results are provided in Appendix.
Figure 6: Generalization test in MVTecAD-loco dataset. We have conducted experiments on detection and segmentation performance in MVTecAD-loco dataset [4] using Patchcore as baseline in one-shot setting. Our method improves the performance for all cases.
@@ 5 Experimental Result
Baselines and Datasets. To compare the effectiveness of the model, we perform a comparative analysis for different baselines, Patchcore [19], Cflow [12], EfficientAD [2], Reverse Distillation [7], using MVTecAD [3], MVTecAD-loco [4], and BTAD [16] datasets. The MVTecAD (MVTec Anomaly Detection) [3] dataset contains over 5,000 high-resolution images divided into fifteen object and texture categories. The BTAD (beanTech Anomaly Detection) [16] dataset contains 2,830 real-world images of 3 industrial products containing body and surface defects. The MVTecAD-loco [4] dataset includes structural and logical anomalies, containing 3,644 images from five categories of real-world industrial inspection scenarios.
Implementation Details. For every baseline, we employ their original setting, and the default backbone network is set to the pre-trained ResNet-18 model, while the reverse-distillation model is a Wide-ResNet50-2. The variance-aware image generator is initialized by the weight parameters of VQGAN [8] pre-trained by ImageNet, and we utilize the pre-trained(ViT-B/16) CLIP model [17]. We set generating 20 images at every iteration. VRAM usage ranges from 2,031MB to 23,277MB and the training time takes 182.6s at every iteration on average. Detailed hyperparameters and explanations of evaluation metrics are provided in the Appendix.
@@ 5.1 Quantitative Results
Compared to the baselines, we have conducted a comparative analysis on the generalization performance of our newly designed framework for one-shot, few shot (5 images), and full-shot (entire training images). Firstly, the integrated average performance of the classes is derived by applying the state-of-the-art model [2] and the latest baselines [19, 12, 7] to the MVTecAD dataset [3]. Also, we have analyzed the five classes with the highest improvement to show the strengths of our method. Secondly, we have conducted experiments using the BTAD dataset [16] consisting of two types of objects and textures, to find the object types in which our model performs well. Finally, we have conducted a final validation of the performance of the model by assigning tasks with complicated object types of the MVTecAD-loco dataset [4].
In Fig. 4, the experimental results on generalized detection using MVTecAD dataset indicate enhanced performance across all baselines, thus verifying our model’s ability to generalize. The results show an average increase of 16.9% in one-shot, 14.3% in few-shot, and 7.8% in full-shot, confirming that our framework can successfully achieve impressive performance even with a few training images. Additionally, we have analyzed the performance of the five highest improving classes in one-shot tasks, as shown in the second and third rows of Fig. 4, respectively. From the results, we can find that object types, such as the cable and metal-nut classes from Patchcore model, show impressive improvements of 19.1% and 18.5%. Other baselines show a similar trend in object types, and the reverse distillation model exhibits a significant increase of 73.3% even in texture types such as leather.
As shown in Fig. 5, we have conducted additional experiments on the BTAD dataset to revalidate which types of defects our model works well. As in the graph of Fig. 5(Top), the detection AUROC shows improvements of 20.2% for object type (1st class) and 4.0% for texture type (2nd class), compared to baseline. Similarly, the segmentation AUROC has also shown greater improvements for the object type, with 4.4% for object types and 1.3% for texture types. In the qualitative results of Fig. 5(Bottom), the object type has resulted in better outcomes for locating the anomaly segments close to the ground truth. As a result, this experiment confirms that our model is stronger on the object types.
However, in the real world, failures happen in more diverse and complicated situations, and thus it is necessary to verify the robustness of our model in more complex settings. MVTec-loco dataset is suitable for testing complex tasks, since it consists of complicated classes, such as logically structured arrangement (logical type) and surface defect recognition (structural type). Fig. 6 shows the experimental results on the MVTec-loco dataset. For the logical type, our method shows improvements of 15.5% for the detection AUROC and 15.8% for segmentation AUROC. Furthermore, for the structural type, our method also shows improvements for both cases, with 10.6% for detection AUROC and 4.3% for segmentation AUROC. In Fig. 6(Bottom), the qualitative results also show that our method can effectively detect and classify the changes in the placement structure of a breakfast box and the subtle differences in a juice bottle.
@@ 5.2 Experimental Analysis
We perform the experimental analysis to validate the effectiveness of our proposed modules and discover the additional intuition for adding generated images. Additional experimental analyses are provided in the Appendix.
@@ 5.2.1 Optimizing results between Our prompt and generated images
The Fig. 7(Top-Left) is a graph comparing the distance based on the similarity between input text and prompts based on the value of cosine similarity. The worst prompt (blue dot) and our prompt (red dot) show a stark difference. The Fig. 7(Top-Right) is a graph comparing the similarity scores of input images and each prompt, and you can see that the prompt received the highest score (33%) and the worst prompt received the lowest score (17%). We show that our prompts have characteristics close to the non-defective image and are evidence that they can boost the performance of text-based knowledge integration modules. The Fig. 7(Bottom), we show that the qualitative image result generated by the best prompt is most similar to the original image. In Table 2, we additionally, quantitative results were confirmed in Fig. 7(Bottom). for single-image evaluation, We compared IS, SSIM, PSNR, and LPIPS scores and showed that our method had the highest diversity and high visual similarity to the original compared to other prompts.

### v2

@@ 2 Related Works
Text-to-Image Generation. Recently, text-to-image generation models [19, 16] are widely studied, which predict or generate the next pixel values based on the given text and all previous predictions of pixel values. For example, DALL-E [20] takes text input from the GPT series model to predict the next pixels, which are then added to the initial image to be placed as the subsequent input along with the text. By repeating the process multiple times, DALL-E can generate an entire image. GAN [12] is known for high-quality and realistic image generation without a prompt, and Contrastive Language-Image Pre-training (CLIP) is another neural network that can determine how well a caption or prompt matches an image. Through VQGAN-CLIP [5], where VQGAN [10] creates an image, CLIP [19] checks whether the generated image from VQGAN matches the text. When the process is repeated, the created image gradually resembles the input text.
@@ 3 Preliminary Analysis
Hypothesis. Data issues related to sensitivity, especially in appearance and level of defectiveness, can vary significantly due to the type of machine used and the capturing conditions. Thus, the similarity between the generated and original images is highly correlated with performance enhancement, and the visual variance of generated images also contributes to performance increases. In Fig. 2, the similarity between the generated and original images is highly correlated with the performance enhancement, and the variance of generated images also contributes to the performance increase. These results empirically validate our hypothesis that the generated images should be similar to the appearance of the provided non-defective images while preserving their visual variance. Furthermore, we compared the correlation between the original image and the naive prompts (0.39), just words (0.35), and ours(0.54), which confirms the necessity of well-designed prompts in our hypothesis.
Figure 2: Correlation of hypothesis. We repeat the tests with different images to confirm our hypothesis. Performance enhancement is strongly connected with the similarity between the generated and original images, and the visual variance of created images also improves performance.
Preliminary Analysis. We repeat the preliminary experiments using various images with a simple scenario to verify our hypothesis in general. In the scenario, we compare the performance between the baseline anomaly detection model and the same model using one additional training image searched on the web or generated by an image generator. We consider a single hazelnut image from the MVTecAD dataset [3], utilizing Patchcore [21] for the baseline anomaly detection model.
@@
Table 1: Preliminary experiment using Patchcore. [21] The results the need for a well-designed generation model to boost anomaly detection performance by adding extra training images.
@@
The results of preliminary experiments are represented in Table 1, showing that performance is affected depending on the generated images. Interestingly, preliminary results show that good image quality does not necessarily help improve performance as presented in the low performance of DALL-E. Meanwhile, given the performance of web images is relatively high, it can be assumed that performance can be improved if elements such as outlines and angles of the original image are preserved in the additional images.
Figure 4: Overview of our framework. Our framework comprises a keyword-to-prompt generator, a variance-aware image generator, and a text-guided knowledge integrator. The keyword-to-prompts generator creates prompts from key input words and selects the best one that matches an input image. A variance-aware image generator creates non-defective images, encoding their visual features into a normal distribution to maintain variance. Our process updates through iteration, and a text-guided knowledge integrator selects the optimal images by comparing the similarity of their latent distribution to the text prompts.
@@ 4 Methodology
As shown in Fig. 4, our framework comprises the keyword-to-prompt generator, variance-aware image generator, and text-guided knowledge integrator. The keyword-to-prompt generator produces a group of prompts consisting of different word combinations based on the essential words presented in the input text, then selects the best prompt that contains comparable information to the input image among the generated ones. The variance-aware image generator generates a set of non-defective images, encoding the visual features of the non-defective images into the normal distribution to preserve their variance. The generation process is iterated with the generator update, and the text-guided knowledge integrator determines the optimal set of generated images by assessing the latent distribution similarity between the image set and the text prompts generated by the keyword-to-prompt generator. Finally, the optimal set of generated images is utilized as the additional training set for the baseline anomaly detection model.
@@ 4.2 Text-guided Knowledge Integrator
Figure 5: Generalization test for anomaly detection in MVTecAD dataset. The first row shows the average improving score across different baselines and varying numbers of non-defective images. The second rows present the average score for the highest-improving five classes.
Figure 7: Generalization test in MVTec-LOCO AD dataset. We have conducted experiments on detection and segmentation performance in MVTec-LOCO AD dataset [4] using Patchcore as baseline in one-shot setting. Our method improves the performance for overall cases.
@@ 5.1 Quantitative Results
We have conducted a comparative analysis of the generalization performance of our newly designed framework for one-shot, few-shot (5 images), and full-shot (entire training images) compared to the baselines. Firstly, the integrated average performance of the classes is derived by applying the state-of-the-art model [2] and the latest baselines [21, 14, 9] to the MVTecAD dataset [3]. Also, we have analyzed the five classes with the highest improvement to show the strengths of our method. Secondly, we have conducted experiments using the BTAD dataset [18] consisting of two types of objects and textures to find the object types in which our model performs well. Finally, we have validated the model’s performance by assigning tasks with complicated object types of the MVTec-LOCO AD dataset [4].
In Fig. 5, the experimental results on generalized detection using MVTecAD dataset indicate enhanced performance across all baselines, thus verifying our model’s ability to generalize. The results show an average increase of 15.7% in one-shot, 14.1% in few-shot, and 5.6% in full-shot, confirming that our framework can successfully achieve impressive performance even with a few training images. Additionally, we have analyzed the performance of the five highest improving classes in one-shot tasks, as shown in the second rows of Fig. 5. The results show that object types, such as the cable and metal-nut classes from the Patchcore model, show impressive improvements of 19.1% and 18.5%, respectively. Other baselines show a similar trend in object types, and the reverse distillation model exhibits a significant increase of 73.3% even in texture types such as leather.
As shown in Fig. 6, we have conducted additional experiments on the BTAD dataset to revalidate which defects our model works well. As in the graph of Fig. 6 (Top), the detection AUROC shows improvements of 20.2% for object type (1st class) and 4.0% for texture type (2nd class), compared to baseline. Similarly, the segmentation AUROC has also shown more significant improvements for the object type, with 4.4% for object types and 1.3% for texture types. In the qualitative results of Fig. 6 (Bottom), the object type has resulted in better outcomes for locating the anomaly segments close to the ground truth. As a result, this experiment confirms that our model is more robust on the object types.
@@
However, failures happen in more diverse and complicated situations in the real world; thus, it is necessary to verify the robustness of our model in more complex settings. MVTec-LOCO AD dataset is suitable for testing complex tasks since it consists of complicated classes, such as logically structured arrangement (logical type) and surface defect recognition (structural type). Fig. 7 shows the experimental results on the MVTec-LOCO AD dataset. For the logical type, our method shows improvements of 15.5% for the detection AUROC and 15.8% for segmentation AUROC. Furthermore, for the structural type, our method also shows improvements for both cases, with 10.6% for detection AUROC and 4.3% for segmentation AUROC. In Fig. 7(Bottom), the qualitative results also show that our method can effectively detect and classify the changes in the placement structure of a breakfast box and the subtle differences in a juice bottle.
@@ 5.2.1 Optimizing results between our prompt and generated images
Fig. 8 (Top-Left) is a graph comparing the distance based on the similarity between input text and prompts based on the cosine similarity value. The worst prompt (blue dot) and our prompt (red dot) show a stark difference. Fig. 8 (Top-Right) is a graph comparing the similarity scores of input images and each prompt, and we can show that the prompt received the highest score (33%) and the worst prompt received the lowest score (17%). We find that our prompts have characteristics close to the non-defective image and are evidence that they can boost the performance of text-based knowledge integration modules. In Fig. 8 (Bottom), we show that the qualitative image result generated by the best prompt is most similar to the original image. In Table 2, we additionally compared the quantitative results based on the qualitative result images in Fig. 8. We compared IS, SSIM, PSNR, and LPIPS scores and showed that our method had the highest diversity and high visual similarity to the original compared to other prompts.
@@
5.2.2 Ablation Studies
Table 3 shows the impact of the respective component of our method. The results with our prompt generator show 4.6% improvement compared to the baseline.
We also analyze the effect of the quantity of the text-based generation images on the performance of the text generator. The model’s performance continuously improves as the number of text-based generation images rises and tends to reach a saturation point. The evidence is that excessive accumulation of non-defective data can act as critical noise in data expression. We found a performance difference of up to 1.1% depending on the number of images.
Lastly, we compare the results of using the variance-aware method. The variance method had an impact of approximately 4.3%. Additionally, In Fig. 10, we compared qualitative results for the Variance-aware method. The Variance-aware method made latent vector from the distribution of $\mathcal{N}\{E_{d},\Sigma_{d}\}$ value in a form close to the normal distribution, and the generation quality is also improved.
6 Conclusion
Acknowledgements: This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (2021-0-01341, Artificial Intelligence Graduate School Program(Chung-Ang University) and 2021-0-02067, Next Generation AI for Multi-purpose Video Search and 2014-3-00123, Development of High Performance Visual BigData Discovery Platform for Large-Scale Realtime Data Analysis) and a grant (22193MFDS471) from the Ministry of Food and Drug Safety in 2024.
Supplementary Material
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 11: Comparison of qualitative results between baseline(Patchcore) and Ours in MVTecAD dataset. Using our model, we show better detection and segmentation performance for each class compared to the baseline.
Appendix A Appendix
Here, we present additional findings that, due to space limitations, could not be included in the main paper. First, detailed implementation, including evaluation metrics and baselines, benchmark datasets, were provided. In addition, the comprehensive quantitative result (Table 5), overall classes performance results by baseline (Tables 6 $\sim$  15) and qualitative experimental results (Figs. 11,  13 $\sim$  14). Additional analysis results and discussion were also provided.
A.1 Comparison of qualitative results in MVTecAD dataset.
In Fig. 11, we can visually confirm that our method is improved over the baseline through the qualitative results of few classes. In Table 4 are few examples of prompts that were finally selected and applied.